OpenStudy (anonymous):

Solve the equation. 9d + 2d + 3d – 4 = 5d

OpenStudy (anonymous):

Add the like terms, so: 9d+2d+3d= 14d. So, 14d-4=5d. You want to get the variables all on one side, so it comes out to be: 9d=4. d=9/4